OpenStudy (anonymous):

Given that the density of alcohol (CH3CH2OH) is 0.789g/mL, convert the mass of alcohol into a volume in a litre of vodka

OpenStudy (anonymous):

\[.789 \frac{g}{mL}(\frac{mL}{.001 L})=789 \frac{g}{L} \]

OpenStudy (anonymous):

how does that work out for you?